OpenStudy (anonymous):

These three points are collinear. (3, 6), (-2, -9), (0, -4) True False

OpenStudy (anonymous):

Collinear points are points that lie on the same straight line. To see if these points lie on the same straight line, see if the slopes of any two points are equal. Use the formula to find the slope between two points, and see if all 3 points are collinear.

OpenStudy (anonymous):

what do you mean?

OpenStudy (amistre64):

the slope formula is:\[\frac{y_o-y_1}{x_o-x_1}\]

OpenStudy (anonymous):

the lines dont match so i guess its false?

OpenStudy (amistre64):

if they dont match, its false

OpenStudy (anonymous):

ok thanks

OpenStudy (amistre64):

my method would be to subtract one point from the other two, and compare (3, 6), (-2, -9), (0, -4) 0, 4 0 , 4 0 ,4 -------------------- 3,10 -2,-5 0,0 10/3 not = 5/2